# Educating LLMs like Human Students: Structure-aware Injection of Domain Knowledge

### v1

@@ 5 Conclusion
This work pioneers in incorporating structure-aware methodologies to enhance domain knowledge injection into large language models (LLMs). Through a novel two-stage training strategy, combining the SCPT and SSFT techniques, we have set a new precedent for efficiently adapting LLMs to specialized domains. The promising results and the scalable data construction framework we developed underscore the viability and potential of our method. The limitations and broader impacts are discussed in Appendix C. We hope to inspire further research in efficient and effective domain adaptation, moving a step closer to models that can truly emulate human intelligence.
@@ A.1 Detailed Setup on LongBench
To focus on the investigation of knowledge injection, we choose 7 subsets from LongBench [4] across single- and multi-document QA tasks in English and Chinese, and the remaining synthetic or code-orientated tasks are eliminated:
  • Single-Doc QA. For single-document QA, we take three subsets from LongBench: (1) Qasper [10], featured by question-answering over NLP technical papers and annotated by NLP practitioners; (2) MultiFieldQA [4], manually curated from multiple data sources and annotated by Ph.D. students; and (3) MultiFieldQA-zh, the Chinese split also provided by Bai et al. [4], covering multiple Chinese scenarios. MultiFieldQA contains 150 Context-Question-Answer triplets to test, and the others adopted subsets include 200 pieces of test samples respectively.
  • Multi-Doc QA. Multi-document QA requires LLMs to extract and combine information from multiple documents to derive the answer, which is generally more challenging than single-doc QA. We take four multi-hop QA datasets: (1) HotpotQA [51], containing 2-hop questions written by native speakers given two related paragraphs; (2) 2WikiMultihopQA [17], involving up to 5-hop questions synthesized through manually designed templates on Wikipedia passages; (3) MuSiQue [41], carefully composed with up to 4-hop reasoning on an increased number of supporting and distracting context evidence; and (4) Dureader [16], developed based on Baidu Search and Baidu Zhidao and filtered by Bai et al. [4] to reduce the data noise. Each subset has 200 test samples.
@@ A.2 Detailed Setup on MMedBench
Data for Evaluation. The Multilingual Medical Benchmark (MMedBench) [35] represents a comprehensive and diverse multilingual medical Question and Answering (QA) benchmark designed to evaluate models’ capabilities of understanding and processing medical content.
@@
Data for Continual Pre-Training. To investigate high-quality domain knowledge injection for LLMs, we collect 18 English textbooks and 33 Chinese textbooks from the National Medical Board Examination in the USA and Mainland China, respectively [21]. All collected textbooks are originally in PDF format and Jin et al. [21] converted them into digital text via OCR and performed some clean-up pre-processing strategies to reduce the data noise. The English and Chinese textbooks count for around 26.1M and 21.5M tokens by Llama2 tokenizer [40]. Then, we randomly sample an extra 28M textbook corpora from MMedC [35] for the other languages (except the unavailable Japanese textbooks), and the final statistics are displayed in Tab. A2.
Table A2: Sample statistics on training data.
@@
Data for Supervised Fine-Tuning. As introduced in Sec. 3.3, we prompt Llama3-70B [2] to build the structure-aware answer explanations on top of the raw SFT samples in MMedBench’s training split, and generate extra QA pairs by traversing the extracted knowledge structure from textbooks. The final quantity statistics are presented in Tab. A2.
Appendix B Resource Requirement
@@
Appendix C Limitations and Discussions
Limitations. A noteworthy limitation of our work, despite its significant achievements, is centered around the exploration of the full potential of our StructTuning strategy for knowledge injection. While our method has successfully demonstrated a remarkable 50% improvement in domain-specific knowledge application and retention using merely 0.3% of the training corpus, the extent of data scalability needed to achieve 100% of the desired effectiveness remains unexplored. This limitation primarily arises from constraints in time and resources available during our research phase. We will delve into the investigations in our future work.
Broader Impacts. We discuss the positive and negative societal impacts as follows:
  • Positive Societal Impacts. Our StructTuning strategy could democratize access to advanced AI, enabling specialized AI assistants across various sectors such as healthcare and education. This method’s efficiency could improve diagnostic tools, personalized learning experiences, and enhanced decision-making processes. Furthermore, by mirroring the human learning process, our approach facilitates the development of AI systems that are easier for people to trust and interact with.
  • Negative Societal Impacts. The enhanced capabilities of LLMs might be exploited for generating convincing yet fraudulent content, raising concerns over misinformation and intellectual deceit. Additionally, if the structured knowledge base contains biases, it could inadvertently reinforce societal stereotypes, necessitating rigorous curation to ensure fairness and accuracy. To counteract potential misuse, controlled model releases and continuous monitoring for abuse are recommended. Ensuring the diversity and contemporaneity of training data can mitigate bias. Collaborative efforts with domain experts, ethicists, and policymakers are vital in developing guidelines for the technology’s ethical use.
Appendix D Prompt Template for Structurization
Fig. A1 displays the few-shot prompt template to query giant commercial LLMs to execute structurization. We first introduce the instruction for structurization (including the task definition and output formats), then provide two representative examples to further illustrate the process. The first example in Fig. A2 tells the model to focus on the existing numerical or enumeration indicators to assist in constructing the aspect-description structure. The second example in Fig. A3 teaches the model to automatically summarize the aspects and attach their corresponding descriptions from the raw text sequences. Finally, in Fig. A1 we ask commercial LLMs to structurize the $input\_statement$ from user input as expected.
Figure A1: Prompt template for few-shot structurization.
Figure A2: The first example for few-shot structurization.
Figure A3: The second example for few-shot structurization.

### v2

@@ 5 Conclusion
This work pioneers in incorporating structure-aware methodologies to enhance domain knowledge injection into large language models. Through a novel SCPT-SSFT paradigm, we have set a new precedent for adapting LLMs to specialized domains, and the promising and scalable results underscore the viability and potential of our method. We hope to inspire further research in efficient and effective domain adaptation, moving a step closer to models that can truly emulate human intelligence.
Limitation. A noteworthy limitation of our work is the under-explored full potential of our StructTuning strategy for knowledge injection. Although we have estimated a scaling law in Eq. 3, the data requirement to achieve 100% of the desired effectiveness remains unverified. Besides, our method introduces negligible data preprocessing costs to extract the domain knowledge structure. However, Sec. B.2 indicates our method can still reduce the overall adaptation cost to 10% since only a small part of training corpus is required. We will delve into the investigations in our future work.
@@ A.1 Detailed Setup on LongBench
To focus on the investigation of knowledge injection, we choose 7 subsets from LongBench (Bai et al., 2023) across single- and multi-document QA tasks in English and Chinese, and the remaining synthetic or code-orientated tasks are eliminated:
  • Single-Doc QA. For single-document QA, we take three subsets from LongBench: (1) Qasper (Dasigi et al., 2021), featured by question-answering over NLP technical papers and annotated by NLP practitioners; (2) MultiFieldQA (Bai et al., 2023), manually curated from multiple data sources and annotated by Ph.D. students; and (3) MultiFieldQA-zh, the Chinese split also provided by Bai et al. (2023), covering multiple Chinese scenarios. MultiFieldQA contains 150 Context-Question-Answer triplets to test, and the others adopted subsets include 200 pieces of test samples respectively.
  • Multi-Doc QA. Multi-document QA requires LLMs to extract and combine information from multiple documents to derive the answer, which is generally more challenging than single-doc QA. We take four multi-hop QA datasets: (1) HotpotQA (Yang et al., 2018), containing 2-hop questions written by native speakers given two related paragraphs; (2) 2WikiMultihopQA (Ho et al., 2020), involving up to 5-hop questions synthesized through manually designed templates on Wikipedia passages; (3) MuSiQue (Trivedi et al., 2022), carefully composed with up to 4-hop reasoning on an increased number of supporting and distracting context evidence; and (4) Dureader (He et al., 2017), developed based on Baidu Search and Baidu Zhidao and filtered by Bai et al. (2023) to reduce the data noise. Each subset has 200 test samples.
For data entries in Single-Doc QA, we extract the knowledge structure for each single passage; in Multi-Doc QA, we identify the knowledge structure across multiple passages for each test sample. There are ultimate 1350 question-answer-passage(s)-(knowledge)structure quadruples to evaluate knowledge injection approaches on LongBench.
@@ A.2 Detailed Setup on MMedBench
Data for Evaluation. The Multilingual Medical Benchmark (MMedBench) (Qiu et al., 2024) represents a comprehensive and diverse multilingual medical Question and Answering (QA) benchmark designed to evaluate models’ capabilities of understanding and processing medical content.
@@
Data for Continual Pre-Training. To investigate high-quality domain knowledge injection for LLMs, we collect 18 English textbooks and 33 Chinese textbooks from the National Medical Board Examination in the USA and Mainland China, respectively (Jin et al., 2020). All collected textbooks are originally in PDF format and Jin et al. (2020) converted them into digital text via OCR and performed some clean-up pre-processing strategies to reduce the data noise. The English and Chinese textbooks count for around 26.1M and 21.5M tokens by Llama2 tokenizer (Touvron et al., 2023b). Then, we randomly sample an extra 28M textbook corpora from MMedC (Qiu et al., 2024) for the other languages (except the unavailable Japanese textbooks), and the statistics are displayed in Tab. A2.
Table A2: Sample statistics on training data. $\dagger$” and $\ddagger$” refer to different sizes.
@@
[…]
Data for Supervised Fine-Tuning. As introduced in Sec. 3.3, we prompt Llama3-70B (Dubey et al., 2024) to build the structure-aware answer explanations on top of the raw SFT samples in MMedBench’s training split, and generate extra QA pairs by traversing the extracted knowledge structure from textbooks. The final quantity statistics are presented in Tab. A2.
Knowledge Structures. We extract the domain knowledge structure for each textbook, where Fig. 3 presents an example, and combine the medical knowledge for six languages in MMedBench. As the (S)CPT corpus for Japanese is collected from Wikipedia rather than textbooks, we derive a single knowledge structure for Japanese medicine.
Figure A1: Knowledge structure extraction from (a) sequential chunks (e.g., from textbooks) by our specialized 7B model and (b) separated trunks (e.g., from websites) by clustering-based methods (Sarthi et al., 2024). Here the terms of “Section”, “Chapter”, and “book” are just examples to help illustrate the knowledge structure.
A.3 Resource Requirement
@@
Appendix B Additional Experiments
B.1 Ablation on Knowledge Structure Extraction
Extracting domain knowledge structure is a prerequisite for subsequent knowledge injection (including both SCPT and SSFT) for language models. In Sec. 3.1, we propose a bottom-up strategy to re-chunk the texts from domain textbooks, summarize a title for each chunk, and send the title list to a specialized 7B model to derive the knowledge structure. The prompt template is displayed in Appendix C.
In fact, we argue that due to the language diversity, a perfectly recovered table of contents of textbooks is unnecessary for domain knowledge injection. A reasonable knowledge structure is sufficient enough. In Tab. A3, we individually adopt few-shot GPT-3.5-Turbo (Brown et al., 2020) and LLaMA3-70B (Dubey et al., 2024) models to extract medical knowledge structure from 18 English textbooks (Jin et al., 2020) (with 26.1M tokens) for subsequent knowledge injection (the backbone LLM is LLaMA2-7B (Touvron et al., 2023b)). Although they present 3.6%-3.7% enhancement on MMedBench (Qiu et al., 2024)’s English test set (denoted as “improvement”), leveraging GPT-3.5-Turbo and LLaMA3-70B is either expensive or time-consuming. GPT-3.5-Turbo costs around 15 dollars to process 26M tokens, while LLaMA3-70B takes around 1.5 hours on 2 A100-80G GPUs, of which both limit the scalability of data pre-processing for structure-aware knowledge injection.
Table A3: Comparison of models to extract knowledge structures on 26M English corpus.
Model Improvement Time Consumption Extra Cost
GPT-3.5-Turbo +3.58 - 15\$
LLaMA3-70B +3.72 1.5h -
Ours-7B +3.69 0.2h -
Inspired by Liu et al. (2024a), we distilled the knowledge structure extraction capability from giant LLMs to a LLaMA2-7B model via supervised fine-tuning. In particular, we instruct LLaMA3-70B to generate 22K training examples (pairs of raw knowledge points and extracted knowledge structures) from Wikipedia, and train a LLaMA2-7B model at a batch size of 128 and a learning rate of 2e-5 for 1 epoch. After utilizing the specialized 7B model to identify the knowledge structure in medical textbooks, as shown in Tab. A3, the results translate to comparable performance on structure-aware knowledge injection. Meanwhile, the inference cost significantly decreases to 0.3 hours, which is more scalable to handle a larger amount of domain corpus.
B.2 Comparison on Training Costs
In Tab. A4, we quantify the total training cost on 8 A100-80G GPUs. According to Qiu et al. (2024), the conventional CPT+SFT paradigm on 25.5B medicine corpus takes more than 30 days to derive the SOTA MMedLM model. In our SCPT+SSFT framework, although the pre-processing (i.e., knowledge structure extraction) introduces an extra 0.6 hours to process 0.3% data (around 76M tokens), the total training process only costs 4.5 hours. As suggested in Fig. 8, when 5% training data is leveraged for knowledge injection to achieve 100% improvement, the overall cost is limited to 3 days, much less than the CPT+SFT approach with more than a month. Those analyses further demonstrate the efficacy and efficiency of our structure-aware knowledge injection framework.
Table A4: Comparison of training costs for knowledge injection.
Paradigm Corpus Improvement Pre-process Total
CPT+SFT 100% 100% - $>$30d
SCPT+SSFT 0.3% 50% 0.6h 4.5h
SCPT+SSFT 5% 100% 9.7h 3d
B.3 F1-Score Evaluation on LongBench
In Sec. 4.1, we mainly follow Zhu & Li (2023b) to investigate the memorization and understanding of injected knowledge by calculating the knowledge recall in models’ responses. Here we report the F1-score measure over the Open-Book QA (OBQA) and Closed-Book QA (CBQA) settings for a thorough comparison. Note that here we use the vanilla question prompt to obtain concise answers, instead of the CoT prompt used in Sec. 4.1 to elicit models’ memorized knowledge. The evaluated models are the same as Sec. 4.1.
In Tab. A5, we report the Open-Book QA (OBQA) baseline for Llama2-7B with passages as inputs, which shows the best performance on MultiFieldQA (MFQA) (Bai et al., 2023) and 2WikiMultihopQA (2Wiki) (Ho et al., 2020) subsets. Then, we establish the Closed-Book QA (CBQA) baseline by traditional CPT+SFT to inject passage contents into model parameters, and supplement the experiment of our SCPT+SSFT technique for comparison. According to the results shown in Tab. A5, CPT+SFT slightly improves the QA performance on several subsets (such as MultiFieldQA-zh (MFQAzh) (Bai et al., 2023)), while the overall F1-Score measure is still inferior to the OBQA performance. In contrast, our SCPT+SSFT approach successfully boosts the closed-book QA performance to 20.1% on average, even surpassing the open-book QA baseline of 18.7%.
The results are consistent with Sec. 4.1, which jointly demonstrate the effectiveness of structure-aware knowledge injection for large language models.
Table A5: F1 Score evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench (Bai et al., 2023) dataset. The best results are marked in bold, and the secondary results are marked with underlines. The backbone model is Llama2-7B (Touvron et al., 2023b).
Task Adaptation SingleDoc-QA MultiDoc-QA Average
Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
OBQA - 19.2 36.8 11.9 25.4 32.8 9.4 5.2 18.7
CBQA CPT+SFT 16.8 23.1 13.2 21.3 19.1 10.4 13.4 16.8
SCPT+SFT 15.2 21.5 15.2 14.9 19.8 5.6 14.1 15.2
SCPT+SSFT 19.7 23.5 19.5 26.4 24.1 12.2 15.4 20.1
B.4 Ablation on SFT Data Synthesis
In Sec. 4.2, we compared our structure-aware knowledge injection with conventional CPT+SFT paradigm on MMedBench. On its English subset, we ablated the training components of our method, and found that the newly synthesized 8K SSFT data (by traversing the extracted knowledge structure) can inspire LLMs’ cross-language capability to apply the learned structured knowledge to solve practical diagnosis problems. Here, we follow Liu et al. (2024b) to randomly generate another 8K QA pairs for SFT alignment for further comparison, denoted as “SFT*”. We randomly sample medical texts and instruct Llama3-70B (Dubey et al., 2024) for data synthesis, without the knowledge structure provided. Tab. A6 indicates that “SFT*” brings slight enhancement to the English test subset, but the average accuracy drops to 34.51% instead. The results further demonstrate our method’s efficacy in the application of the injected, structured domain knowledge.
Table A6: Comparison of SFT data synthesis strategies on MMedBench. The backbone LLM is the same Llama2-7B model after SCPT on English textbooks.
SFT synthesis English Chinese Japanese French Russian Spanish Average
- 46.50 32.14 20.10 18.17 53.91 39.97 35.13
SFT* 47.13 32.49 16.58 16.72 51.95 42.16 34.51
SSFT* 49.10 33.92 18.33 27.14 57.42 43.73 38.27
Appendix C Prompt Template for Knowledge Structure Extraction
Fig. A2 displays the prompt template to query our specialized 7B model to extract knowledge structure on given knowledge points, which introduces the task definition, detailed instruction, and output formats to illustrate the process.
Figure A2: Prompt template for knowledge structure identification.